OpenStudy (anonymous):

if there is some food for 500 men for 6 days, after 2 days 100 people departure that place then food will be sufficient for how many days?

OpenStudy (anonymous):

500*6=3.000 500*2= 1.000 3.000-1.000= 2.000 (500-100)d= 2000 400d=2.000 d=2000/400 d=5days

OpenStudy (anonymous):

what is d here

OpenStudy (anonymous):

d is days

OpenStudy (anonymous):

you also could have done it as a ratio

OpenStudy (anonymous):

now tell me another thing when 500 men are multiplied with 6 days then we get 3000 what is the unit of 3000?

OpenStudy (anonymous):

3000 men/days?

OpenStudy (anonymous):

we can do it with ratio but in case of proportion food must be constant while in que food will be decreased after 2 days?then how can we do it by proportion method

OpenStudy (anonymous):

You are right you cant I read the question incorrectly

OpenStudy (anonymous):

i can not understand the unit of 3000 plz explain it if you can

OpenStudy (anonymous):

500 x 6 mandays - 500 x 2 mandays = days x 400 men Thus, days = 5